OpenStudy (anonymous):

Find the general term set {8,12,16, 24....}

OpenStudy (campbell_st):

well its an arithmetic sequence since you add 4 all the time... so you can use the formula for the general term \[a{_n} = a + (n - 1)\times d\] a = 1st term, n = number of terms and d is the common difference in your question a = 8 and d = 4 just substitute and evaluate